# Perfect codes over non-prime power alphabets: an approach based on Diophantine equations

### v1

@@ Abstract.
The classification of perfect codes over non-prime power alphabets has been an open problem for which there have been no new results in almost $50$ years. In this paper, we show non-existence of perfect $2-$error correcting codes over $q-$ary alphabets for more than 170 new values of $q$. Our methods rely on techniques from the resolution of generalised Ramanujan–Nagell equations and from computational number theory.
@@ 1.1. Background
Error correcting codes have been widely studied since the 1940s, both due to their mathematical interest and for their applicability in the transmission of information over noisy channels. In this paper, we shall consider the following standard definition of an error correcting code.
@@
It is a standard fact in coding theory (for example, see [10, Theorem 2.16]) that any $e-$error correcting code with parameters $(n,M)$ over an alphabet of size $q$ satisfies the sphere packing bound or Hamming bound:
$M\left(\sum_{j=0}^{e}{\binom{n}{j}}(q-1)^{j}\right)\leq q^{n}.$
In the case where the previous expression is an equality, we say that the code is perfect. In other words, $C$ is perfect if and only if $n$ and $M$ satisfy the following Diophantine equation:
@@
The classification of perfect codes was a very active topic of research in the late 20th century, and we refer the reader to [9] for an excellent survey on the most relevant results.
In particular, we highlight that if $q$ is a prime power, the parameters of $q-$ary perfect codes have been completely classified by Tietäväinen [22], and independently by Leontiev and Zinoniev [15].
On the other hand, if $q$ is not a prime power and $e\geq 3$, the work by Reuvers [19], Best [4] and Hong [11] shows that there are no perfect $e-$error correcting codes over any $q-$ary alphabet. In fact, many authors conjecture (see [9, Section 5], for example) the following:
@@
While this question has been completely settled if $e\geq 3$, non-existence results are much more scarce if $e\leq 2$. For instance, if $e=2$, the work of Reuvers [19], van Lint [17] and Bassalygo et al. [2] show that perfect $2-$error correcting codes do not exist for
$q=6,10,15,21,22,26,30,35\quad\text{and}\quad q=2^{s}3^{r},\text{ where }s,r>0.$
However, these results date back to the 1970s, and it is surprising to note that there have been no new results since. In this paper, we address this gap in the literature by proving new non-existence results for the case $e=2$ and 172 new values of $q$, providing further evidence for Conjecture 1.
@@ 1.2. The main results
Since our methodology applies for $q=6,15,21,22,26,30,35$, for $q=10$, and for some cases where $q=2^{r}\cdot 3^{s}$, we shall recover the aforementioned results by Reuvers [19], van Lint [17] and Bassalygo et al [2].
@@
The reason why we need to exclude $q=94$ and $q=166$ from Theorem 2 is due to computational difficulties. We shall comment more on this fact in Section 5.
@@
In order to prove Theorems 2 and 3, we will study the Hamming bound (1) with $e=2$. We shall see that, for a fixed value of $q$, it can be reduced to a Ramanujan–Nagell type equation, which we can solve effectively with the use of several techniques from computational number theory. If, after resolving the equation, there are any solutions $(n,M)$ to (1), we can use Lloyd’s theorem to prove that they cannot constitute a perfect code.
The organisation of this paper is as follows. In Section 2, we will reduce the Hamming bound to a Ramanujan–Nagell equation and prove Theorem 3. In Section 3, we will build upon the work of von Känel and Matschke [12] in order to obtain all solutions to the previously obtained Ramanujan–Nagell equations. In Section 4, we will show that the outstanding solutions cannot form perfect codes by using Lloyd’s theorem, and therefore prove Theorem 2. Finally, in Section 5, we shall discuss complications arising from our methodology, as well as potential future lines of work.
All computations in this paper have been implemented in Magma [5] and are available for the reader in the author’s GitHub repository: https://github.com/PJCazorla/perfect-q-ary-codes.
Acknowledgements: The author would like to thank Yuri Bazlov for insightful discussions.
@@ 2. The Hamming bound and Ramanujan–Nagell type equations
We suppose that there exists a perfect 2 error-correcting $q-$ary code with parameters $(n,M)$. In order to avoid trivial codes, we shall assume that $n\geq 5$ and $M\geq 2$. Then, it follows from the Hamming bound (1) with $e=2$ that
@@
$A=\{q_{1}^{\alpha_{1}}\dots q_{k}^{\alpha_{k}}\mid\alpha_{1},\dots,\alpha_{k}\geq 0\}.$
@@
(6) $x^{2}+b=cy,\quad x\in\mathbb{Z},\quad y\in A,$
where $b=D_{q}$ and $c=8$. This is an example of a generalised Ramanujan–Nagell equation, as considered by von Känel and Matschke in [12]. We can immediately use their work to prove Theorem 3.
@@
By [12, Corollary K], it follows that there are at most finitely many solutions to any generalised Ramanujan–Nagell equation such as (6) and, consequently, there are at most finitely many solutions to (5), showing that there can only be finitely many perfect $2-$error correcting $q-$ary codes.
@@
By the work of Siegel [20], these curves have finitely many integral points. In addition, by (8), there are only finitely many curves $E_{d}$ to consider. Consequently, there are at most finitely many solutions to (6), finishing the proof. ∎
@@
$x^{\prime}=\frac{x}{2}\quad\text{and}\quad D^{\prime}_{q}=\frac{D_{q}}{2}.$
@@
$(x^{\prime})^{2}+(D^{\prime}_{q})=2q_{1}^{n_{1}}\dots q_{k}^{n_{k}}.$
@@ 3. Resolving the Diophantine equations
As we explained in Section 2, the existence of a perfect $2-$error correcting code over a $q-$ary alphabet implies the existence of a solution $(x,y)$ to the generalised Ramanujan–Nagell equation (6). In this section, we shall present an algorithm to solve this type of equations. Our algorithm is implemented in Magma in the GitHub repository https://github.com/PJCazorla/perfect-q-ary-codes and is based on [12, Algorithm 6.2].
@@
$x=\frac{Y}{d}\quad\text{and}\quad y=\frac{d}{c_{0}}\left(\frac{X}{c_{1}d}\right)^{3}.$
In order to compute the integral points on the Mordell curves $E_{d}$, we distinguish two cases. Following [12, Section 4.1], we define $a_{S}$ by
$a_{S}=1728\prod_{p}p^{\min\{\text{ord}_{p}(bd^{2}),2\}},$
where the product is taken over all primes $p$ and $\text{ord}_{p}(.)$ denotes the standard $p-$adic valuation. If $a_{S}<500,000$, we can use [12, Algorithm 4.2] to compute all integral points very efficiently. This restriction is due to the fact that the algorithm requires the computation of all elliptic curves of conductor $N\mid a_{S}$. If $a_{S}<500,000$, these curves have been computed by Cremona [6] and so the necessary computational effort is minimal.
@@
We also note that if $r=0$, there is no need to compute any generators to the Mordell-Weil group. Similarly, if $r=1$, a generator can be found by computing Heegner points, as shown by Gross and Zagier [8]. We implement these two improvements in our code. While it could seem that these are not significant for the computation, they actually help tremendously and allow us to solve many equations which would otherwise not be amenable to our techniques.
@@ 4. Dealing with the outstanding solutions: Lloyd’s theorem
(Lloyd) Let $n,e$ and $q$ be integers. Suppose that there exists a perfect $e-$error correcting code over a $q-$ary alphabet with word length $n$. Then, the polynomial
$L_{e}(x)=\sum_{i=0}^{e}(-1)^{i}(q-1)^{e-i}\binom{x-1}{i}\binom{n-x}{e-i}$
@@
$L_{2}(x)=(q-1)^{2}\frac{(n-x)(n-x-1)}{2}-(q-1)(x-1)(n-x)+\frac{(x-1)(x-2)}{2}$
has two distinct integer roots $x_{1},x_{2}$ with $1\leq x_{1}<x_{2}\leq n$. However, we check with Magma that this is not the case for any of the values of $q$ and $n$ in Table 1.
@@
Consequently, for all the values of $q$ in the statement of the Theorem, there does not exist a perfect $2-$error correcting code over an alphabet of size $q$. ∎
5. Concluding remarks
We highlight that, with the methodology that we present here, it seems unlikely that the values of $q$ in Theorem 2 can be extended further.
This is due to the fact that, despite the computational improvements that we have introduced in Section 3, determining generators of the group $E_{d}(\mathbb{Q})$, where $E_{d}$ is the Mordell curve (11), is very hard computationally as the prime divisors of $q$ grow. For instance, while trying to solve the Ramanujan–Nagell equation arising from the case $q=94=2\cdot 47$, we need to find all integral points on the Mordell curve
$E_{2\cdot 47^{2}}:y^{2}=x^{3}+161478403652,$

### v2

@@ Abstract.
Perfect error correcting codes allow for an optimal transmission of information while guaranteeing error correction. For this reason, proving their existence has been a classical problem in both pure mathematics and information theory. Indeed, the classification of the parameters of $e-$error correcting perfect codes over $q-$ary alphabets was a very active topic of research in the late 20th century. Consequently, all parameters of perfect $e-$error correcting codes were found if $e\geq 3$, and it was conjectured that no perfect $2-$error correcting codes exist over any $q-$ary alphabet, where $q>3$. In the 1970s, this was proved for $q$ a prime power, for $q=2^{r}3^{s}$ and for only $7$ other values of $q$. Almost $50$ years later, it is surprising to note that there have been no new results in this regard and the classification of $2-$error correcting codes over non-prime power alphabets remains an open problem. In this paper, we use techniques from the resolution of generalised Ramanujan–Nagell equation and from modern computational number theory to show that perfect $2-$error correcting codes do not exist for $172$ new values of $q$ which are not prime powers, substantially increasing the values of $q$ which are now classified. In addition, we prove that, for any fixed value of $q$, there can be at most finitely many perfect $2-$error correcting codes over an alphabet of size $q$.
@@ 1.1. Background
Error correcting codes have been widely studied since the 1940s. Apart from their intrinsic mathematical interest, this is due to the fact that error correcting codes have many useful applications for information transmission and engineering. For example, in [6] the authors discuss possible applications of error-correcting codes to semiconductor memory, while the classical article by Shannon [26] introduced error-correcting codes as a way to transmit information over noisy channels.
While these applications are one of the main reasons why error correcting codes are studied, the developments in this paper will be theoretical. Throughout the article, we shall consider the following standard definition of an error correcting code.
@@
It is a standard fact in coding theory (for example, see [12, Theorem 2.16]) that any $e-$error correcting code with parameters $(n,M)$ over an alphabet of size $q$ satisfies the sphere packing bound or Hamming bound:
(1) $M\left(\sum_{j=0}^{e}{\binom{n}{j}}(q-1)^{j}\right)\leq q^{n}.$
Let us briefly explain why this inequality holds. For this, let us define the set
(2) $B(\overline{w},d)=\{\overline{w}_{2}\in C\mid\overline{w}\text{ and }\overline{w}_{2}\text{ differ in at most }d\text{ positions.}\},$. } ,
where $\overline{w}\in C$ is fixed and $d\geq 0$ is an integer. It is an easy combinatorial exercise to see that
(3) $|B(\overline{w},d)|=\sum_{j=0}^{d}{\binom{n}{j}}(q-1)^{j}.$
In order for a code to be an $e-$error correcting code, it follows that the sets $B(\overline{w},e)$ and $B(\overline{w}^{\prime},e)$ are disjoint for $\overline{w}\neq\overline{w}^{\prime}$, and therefore, we have that
(4) $\left|\bigcup_{\overline{w}\in C}B(\overline{w},e)\right|=M\left(\sum_{j=0}^{d}{\binom{n}{j}}(q-1)^{j}\right).$
Since there are precisely $q^{n}$ elements in $Z_{q}^{n}$, inequality (1) follows. In the case where (1) is an equality, we say that the code is perfect. In other words, $C$ is perfect if and only if $n$ and $M$ satisfy the following Diophantine equation:
@@
Perfect codes are optimal for information transmission, in the sense that they allow for the maximum possible number of words while being able to correct $e$ errors. In other words, perfect codes are the most expressive $e-$error correcting codes that are theoretically possible, and it is desirable to know whether they exist for a fixed value of $q$ and $e$.
For these reasons, the classification of perfect codes was a very active topic of research in the late 20th century, and this led to the production of a vast amount of literature on the topic. The most relevant works are summarised in Table 1 and we refer the reader to [11] for a detailed survey on the history of perfect codes.
Reference $q$ $e$ Main results
Tietäväinen [29], and Leontiev and Zinoniev [18] (independently) Prime power $\geq 1$ All perfect codes over prime power alphabets are either Hamming codes or have parameters $(n,M,$ $e,q)=(11,3^{6},2,3),(23,2^{12},3,2)$.
Reuvers [25] Non- prime power 3, 4, 5 There are no perfect $e-$error correcting codes over a $q$-ary alphabet.
Best [4] 7, $\geq 9$
Hong [13] $6$, $8$
Reuvers [25] 6, 15, 21, 22, 26, 30, 35 2 There are no perfect $e-$error correcting codes over a $q$-ary alphabet.
van Lint [20] 10
Bassalygo et al. [2] $2^{r}\cdot 3^{s}$
Table 1. Summary of existing results on perfect $e-$error correcting codes over $q-$ary alphabets
From Table 1, we immediately notice that perfect codes do exist if $q$ is a prime power, and are either Hamming codes with parameters
(6) $(n,M,e,q)=\left(\frac{q^{r}-1}{q-1},\frac{q^{r}-1}{q-1}-r-1,1,q\right),$
for any $r\geq 2$, or are Golay codes with parameters
(7) $(n,M,e,q)=(11,3^{6},2,3),(23,2^{12},3,2).$
This was proved by Tietäväinen [29], and independently by Leontiev and Zinoniev [18]. Therefore, if $q$ is a prime power, perfect codes do exist and are completely classified for any value of $e\geq 1$.
However, as it is apparent from Table 1, the situation when $q$ is not a prime power is quite different. In this case, all the existing results are negative for any values of $e$ and $q$. This has led many authors to conjecture (see for example [11, Section 5]) the following:
@@
In Table 1, we see that the combination of the works of Reuvers [25], Best [4] and Hong [4] successfully proves Conjecture 1 for $e-$error correcting codes, where $e\geq 3$. For $e\leq 2$, the number of results is much more scarce, and almost non-existent for $e=1$, as remarked by Heden in [11, Section 2].
For $e=2$, as shown in the last rows of Table 1, the existing results are due to Reuvers [25], van Lint [20] and Bassalygo et al. [2]. Together, they show that perfect $2-$error correcting codes do not exist for
(8) $q=6,10,15,21,22,26,30,35\quad\text{and}\quad q=2^{s}3^{r},$
where $r,s>0$ are positive integers. We remark that the methods used in all these papers are specific to the cases at hand and cannot be generalised to other values of $q$. This is probably the reason why there have been no new results in the area in more than $50$ years.
Indeed, we believe that the great difficulty of dealing with the case where $q$ is not a prime power, as well as the lack of a general methodology for these cases justifies the scarcity of literature.
To illustrate this, we note that there is a significant amount of modern articles which aim to generalise the results in Table 1 to extended codes or to different metrics, but they always treat the case when $q$ is a prime power. For example, under this assumption Li and Xing [19] classify perfect quantum codes and Gubitosi, Portela and Qureshi [10] show analogous results for the Niederreiter-Rosenbloom-Tsfasman metric.
In conclusion, we identify a clear gap in the literature in the case $e=2$, since, apart from the family studied in [2], there are only $7$ values of $q$ which are not prime powers and for which non-existence has been proved, and there have been no new results in half a century. In addition, the techniques used in the proofs cannot be generalised, which makes it difficult to extend the results to different codes or metrics.
In this paper, we address this gap in the literature by proving non-existence results for the case $e=2$ and 172 new values of $q$, therefore substantially increasing the evidence towards Conjecture 1. In addition, our methods are theoretically valid for all values of $q$ and the only obstructions that we encounter are of a computational nature. Consequently, it seems feasible to generalise them to other problems, and we shall comment more on this in Section 5.2.
@@ 1.2. The main results
Since our methodology applies for $q=6,15,21,22,26,30,35$, for $q=10$, and for some cases where $q=2^{r}\cdot 3^{s}$, we shall recover the aforementioned results by Reuvers [25], van Lint [20] and Bassalygo et al [2]. Therefore, our findings are consistent with the previously existing literature.
@@
The values of $q$ in the theorem are optimal given the current existing computational techniques and the methods used in this paper. We will comment more on this in Section 5.1.
@@
In order to prove Theorems 2 and 3, we will study the perfect code condition (5) with $e=2$. We shall see that, for a fixed value of $q$, it can be reduced to a Ramanujan–Nagell type equation, which we can solve effectively with the use of several techniques from computational number theory.
The history of Ramanujan–Nagell type equations is rich and has motivated a vast amount of research in the number theory community ever since Ramanujan conjectured that the Diophantine equation
(9) $x^{2}+7=2^{n},$
only had solutions for $n=3,4,5,7$ and $15$. This was proved by Mordell in 1948 [23] and led to many researchers considering generalisations of (9). We refer the reader to [16] for a detailed survey of generalisations of the Ramanujan–Nagell equation.
In order to prove Theorems 2 and 3, we shall solve a Ramanujan–Nagell type equation. If, after resolving the equation, there are any solutions $(n,M)$ to (5), we can use Lloyd’s theorem to prove that they cannot constitute a perfect code.
The organisation of this paper is as follows. In Section 2, we will reduce the Hamming bound to a Ramanujan–Nagell equation and prove Theorem 3. In Section 3, we will build upon the work of von Känel and Matschke [14] in order to obtain all solutions to the previously obtained Ramanujan–Nagell equations. In Section 4, we will show that the outstanding solutions cannot form perfect codes by using Lloyd’s theorem, and therefore prove Theorem 2. In Section 5, we shall discuss complications arising from our methodology, as well as potential future lines of work. Finally, in Section 6, we will briefly summarise our findings and explain how they overcome the shortcomings in the literature.
All computations in this paper have been carried out with the use of Magma [5] code. The code is available for the reader in the author’s GitHub repository (https://github.com/PJCazorla/perfect-q-ary-codes) and can be used to check the correctness of the computations.
@@ 2. The Hamming bound and Ramanujan–Nagell type equations
We suppose that there exists a perfect 2 error-correcting $q-$ary code with parameters $(n,M)$. In order to avoid trivial codes, we shall assume that $n\geq 5$ and $M\geq 2$. Then, it follows from the perfect code condition (5) with $e=2$ that
@@
(14) $A=\{q_{1}^{\alpha_{1}}\dots q_{k}^{\alpha_{k}}\mid\alpha_{1},\dots,\alpha_{k}\geq 0\}.$
@@
(15) $x^{2}+b=cy,$
where $x\in\mathbb{Z}$, $y\in A$, $b=D_{q}$ and $c=8$. This is an example of a generalised Ramanujan–Nagell equation, as considered by von Känel and Matschke in [14]. We can immediately use their work to prove Theorem 3.
@@
By [14, Corollary K], it follows that there are at most finitely many solutions to any generalised Ramanujan–Nagell equation such as (15) and, consequently, there are at most finitely many solutions to (13), showing that there can only be finitely many perfect $2-$error correcting $q-$ary codes for a fixed value of $q$.
@@
By the work of Siegel [27], these curves have finitely many integral points. In addition, since the set in (17) is finite, there are only finitely many curves $E_{d}$ to consider. Consequently, there are at most finitely many solutions to (15), finishing the proof. ∎
@@
(22) $x^{\prime}=\frac{x}{2}\quad\text{and}\quad D^{\prime}_{q}=\frac{D_{q}}{2}.$
@@
(23) $(x^{\prime})^{2}+D^{\prime}_{q}=2q_{1}^{n_{1}}\dots q_{k}^{n_{k}}.$
@@ 3. Resolving the Diophantine equations
As we explained in Section 2, the existence of a perfect $2-$error correcting code over a $q-$ary alphabet implies the existence of a solution $(x,y)$ to the generalised Ramanujan–Nagell equation (15). In this section, we shall present an algorithm to solve this type of equations. Our algorithm is based on [14, Algorithm 6.2].
@@
(24) $x=\frac{Y}{d}\quad\text{and}\quad y=\frac{d}{c_{0}}\left(\frac{X}{c_{1}d}\right)^{3}.$
In order to compute the integral points on the Mordell curves $E_{d}$, we distinguish two cases. Following [14, Section 4.1], we define $a_{S}$ by
(25) $a_{S}=1728\prod_{p}p^{\min\{\text{ord}_{p}(bd^{2}),2\}},$
where the product is taken over all primes $p$ and $\text{ord}_{p}(.)$ denotes the standard $p-$adic valuation. If $a_{S}<500,000$, we can use [14, Algorithm 4.2] to compute all integral points very efficiently. This restriction is due to the fact that the algorithm requires the computation of all elliptic curves of conductor $N\mid a_{S}$. If $a_{S}<500,000$, these curves have been computed by Cremona [7] and so the necessary computational effort is minimal.
@@
We also remark that if $r=0$, there is no need to compute any generators to the Mordell-Weil group. Similarly, if $r=1$, a generator can be found by computing Heegner points, as shown by Gross and Zagier [9]. We implement these two improvements in our code. While it could seem that these are not significant for the computation, they actually help tremendously and allow us to solve many equations which would otherwise not be amenable to our techniques.
@@ 4. Dealing with the outstanding solutions: Lloyd’s theorem
(Lloyd) Let $n,e$ and $q$ be positive integers. Suppose that there exists a perfect $e-$error correcting code over a $q-$ary alphabet with word length $n$. Then, the polynomial
(28) $L_{e}(x)=\sum_{i=0}^{e}(-1)^{i}(q-1)^{e-i}\binom{x-1}{i}\binom{n-x}{e-i}$
@@
(29) $L_{2}(x)=(q-1)^{2}\frac{(n-x)(n-x-1)}{2}-(q-1)(x-1)(n-x)+\frac{(x-1)(x-2)}{2}$
@@
has two distinct integer roots $x_{1},x_{2}$ with $1\leq x_{1}<x_{2}\leq n$. However, we check with Magma that this is not the case for any of the values of $q$ and $n$ in Table 2. Consequently, for all the values of $q$ in the statement of Theorem 2, there does not exist a perfect $2-$error correcting code over an alphabet of size $q$. ∎
5. Limitations and future lines of work
In this section, we shall briefly comment on the limitations of the methodology that we present in this paper, as well as propose future lines of work and different problems which could be solved by leveraging these techniques.
5.1. Limitations of our methodology
Firstly, we remark that our methods can show non-existence of perfect $2-$error correcting codes for a fixed value of $q$, but they are not able to handle infinite families of $q$. The main reason why is due to the fact that the coefficient $D_{q}$ in (13) depends on $q$ and it would no longer be constant.
For this reason, even if the set (17) remains finite (which could be achieved by fixing the prime divisors of $q$, for example), we would effectively need to obtain all integral points for a family of Mordell curves $E_{d(q)}$. In this situation, the results of Siegel [27] do not apply and, consequently, it is entirely possible that there are infinitely many integral points to consider. In addition, and to the best of our knowledge, there are not any techniques that allow to determine integral points over families of elliptic curves in an effective manner. This is why we believe that, in order to prove Conjecture 1, even just for $e=2$, significantly new ideas will need to be introduced.
If we fix $q$, we note that, in principle, all the techniques that we have presented in this paper would be applicable for any $q$. However, in practice, it is computationally unfeasible to resolve (13) for values of $q$ other than those in Theorem 2. For this reason, it is unlikely that the range of values of $q$ can be extended much further with our methodology.
This is mainly due to the fact that, despite the computational improvements that we have introduced in Section 3, we ultimately need to find a list of generators for the group $E_{d}(\mathbb{Q})$, where $E_{d}$ is the Mordell curve (21). As the conductor of $E_{d}$ grows, even finding one rational point on $E_{d}$ is a very hard problem and, consequently, it is impractical to compute a basis for $E_{d}(\mathbb{Q})$.
To illustrate the computational difficulties, let us consider the case $q=94=2\cdot 47$. This is excluded from Theorem 2 because our methodology would require us to work with the curve
(30) $E_{2\cdot 47^{2}}:y^{2}=x^{3}+161478403652,$